OpenStudy (anonymous):

The position of a mass on a spring is given by x = (8.0 cm) cos [2 π t / (0.96 s)]. a) What is the period, T, of this motion? b) Where is the mass at t = 3.9 s?

OpenStudy (unklerhaukus):

\[x(t) = 8 \cos \frac{2 π }{0.96}t\] Compair with\[x(t) = A\cos(ωt + φ)=A\cos(2\pi ft + φ)=A\cos\left(\frac{2\pi }Tt + φ\right)\]where \(A\) is the amplitude \(\omega\) is the angular frequency the frequency is \(f=\frac{\omega}{2\pi}\) the period is \(T=\frac1{f}=\frac{2\pi}\omega\) \( φ \) is the phase lag